The market below exists for electricity produced by coal at a power plant in Tennessee. The marginal private cost (MPC) curve represents the private costs of the power plant to produce electricity from coal. Coal-fired electricity generation, however, also produces an externality. Production creates emissions that can harm human health, contribute to greenhouse gas emissions, and degrade ecosystems. The marginal social cost (MSC) curve represents the cost of coal-fired electricity production in Tennessee, including the costs borne by society. Use the following graph to answers questions 1 - 9:

MSC MPC Тв F1 I- Demand Q Q Quantity of Coal-fired Electricity

What is the market equilibrium price  and quantity ?

What is the socially efficient price  and quantity  ?

At the market equilibrium outcome, what is the total external cost?

At the market equilibrium outcome, what is consumer surplus?

At the market equilibrium outcome, what is the producer surplus?

At the social efficient equilibrium outcome, what is the total external cost?

At the socially efficient outcome, what is consumer surplus?

At the socially efficient outcome, what is producer surplus?

If the price changes from P2 to P1, what is the CHANGE in net benefits to society? Show you work!
Answer #1

A) market equilibrium price at Intersection of MPC and demand curve. So market price is p2 amd Quantity Q2.

B) social efficent QUANTITY is at Intersection of MSC and demand. So social efficent QUANTITY is Q1 and price p1.

C) Total external cost=d+f+g+k+h

D) CONSUMERs surplus=A+B+f+j

E) produer surplus=C+d+h

F) total external cost=d+f

G) CONSUMERs surplus=A

H) produer surplus=C+b+f+d

i)change in net benefit=K